OpenStudy (anonymous):

finish problem: 4pm2sqrt{8}/8

OpenStudy (chaise):

\[\frac{4\pm \sqrt{8}}{8}\] \[=\frac{4\pm 2\sqrt{2*4}}{8}\] \[=\frac{4\pm 4\sqrt{2}}{8}\] \[\approx 1.2\]

OpenStudy (anonymous):

why did u leave it in dec. form?

OpenStudy (chaise):

I didn't have to. I could have left it as: \[\frac{1}{2}+\frac{1}{\sqrt{2}}\] \[\approx 1.2\]

OpenStudy (anonymous):

and this will be -1.2,+1.2?